OpenStudy (anonymous):

If you multiply a number by four and then subtract negative ten, the difference is negative eighteen. What is the number?

OpenStudy (anonymous):

4x-(-10)=18 4x+10=18 4x=8 x=2

OpenStudy (anonymous):

4x-(-10)=18 4x+10=18 4x=8 x=2

OpenStudy (triciaal):

@libbyanne40 "the difference is negative eighteen" can you clarify what difference @shahd-owies you assume the result is 18 and then made an error the original has negative 18

OpenStudy (triciaal):

@shahd-owies sorry that is not an assumption. By definition the difference is the result from subtraction. you only have an error

OpenStudy (anonymous):

you are right @triciaal i am sorry 4x-(-10)=-18 4x+10=-18 4x=-28 x=-7